OpenStudy (anonymous):

It takes 16 pounds of seed to completely plant a 3-acre field. How many pounds of seed are needed per acre?

OpenStudy (anonymous):

\[\frac{ 16 \ pounds }{ 3 \ acres } = \frac{ 5\frac{ 1 }{ 3 } \ pounds }{ 1 \ acre }\]